OpenStudy (anonymous):

simplest form (8x^-6)^1/3

OpenStudy (anonymous):

\[8^{-6^{1/3}} = 8^{-6*1/3} = 8^{-2}\]

OpenStudy (anonymous):

it doesn't look like that

OpenStudy (anonymous):

the parenthesis change the answer

OpenStudy (anonymous):

and there is a X not multiplication but a variable

OpenStudy (anonymous):

\[2\times \sqrt[3]{1/x^{6}}\]

OpenStudy (anonymous):

the possible answers are 8/x^2 , x^2/2 , x^6/ 2, or 2/X^2

OpenStudy (anonymous):

those are fractions Btw

OpenStudy (anonymous):

8^1/3 = 2 , x^-6 = 1/(x^6) , (whatever)^1/3 = \[\sqrt[3]{whatever}\] in this case "whatever" is "1/(x^6)"

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so none of the ones i put up there would work?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

actually my mistake, it could be 2/x^2

OpenStudy (anonymous):

you are right! my answer wasn't simple enough. 2/x^2 is correct answer!!! sorry